OpenStudy (anonymous):

Put your answer in lowest terms, in descending powers of a, and place variables alphabetically: a^2-b^2/a^3-b^3=________

OpenStudy (shubhamsrg):

a^2 - b^2 =(a+b)(a-b) a^3 - b^3 = (a-b)(a^2 -ab + b^2) just make these substituions and you should get your ans

OpenStudy (anonymous):

What do you mean by make these substitutions, do I factorize them and then put it in alphatebical order?

OpenStudy (shubhamsrg):

these are already factorised.. just put the RHS of a^2-b^2 in numerator and RHS of a^3 -b^3 in denominator,,cancel out same terms and voila! you get your ans!

OpenStudy (anonymous):

what does RHS stad for?

OpenStudy (anonymous):

*stand

OpenStudy (shubhamsrg):

right hand side.. like in x=4 x is LHS (left hand side) and 4 is RHS(right hand side)

OpenStudy (anonymous):

Okay and do you know how I can factor x^2+4xy+4y^2?

OpenStudy (shubhamsrg):

can you tell me the expansion of (x+y)^2 ?

OpenStudy (anonymous):

x^2+2xy+y^2 ?

OpenStudy (shubhamsrg):

yep..and what if i ask (x + 2y)^2 ?

OpenStudy (anonymous):

x^2+4xy+2y^2 ?

OpenStudy (shubhamsrg):

bingo!! isnt that what you were looking for?

OpenStudy (shubhamsrg):

ohh..not 2y^2,,its (2y)^2 i.e 4y^2

OpenStudy (anonymous):

Yeah, that is where I get lost..

OpenStudy (shubhamsrg):

so,,you get your answers right?

OpenStudy (anonymous):

Could you factorize all the trinomial please? Because if I do it with the 2y all to the ^2 it does not give me a right answer, it just gets all messed up

OpenStudy (shubhamsrg):

hmmn,,you want only the answer right?

OpenStudy (anonymous):

no, Im saying could you explainn the whole thing

OpenStudy (shubhamsrg):

alright,,leme try in your first ques we have (a+b)(a-b)/ (a-b)(a^2 -ab +b^2) => (a+b)/(a^2 - ab +b^2) --> which i guess should be the ans [note that (a-b) is common and gets cancelled out] in the second ques..you yourself reached the ans.. it is (x+2y)^2..

OpenStudy (anonymous):

Okay, thank you

OpenStudy (shubhamsrg):

hmm..glad to have helped ^_^